Compute for the standard error of the mean given the population variance and the sample size. Show your complete solutions.3.1. 𝜎" = 20.50; 𝑛 = 33.2. 𝜎" = 13.18; 𝑛 = 5 


1
Expert's answer
2021-05-07T09:26:07-0400

3.1 "\sigma''=20.5, n=3"


standard error of the mean "=\sqrt{\sigma'' n}"

"=\sqrt{20.5\times 3}

 \\=\sqrt{61.5}=7.84"


3.2 "\sigma''=13.18, n=5"


standard error of the mean "=\sqrt{\sigma'' n}"

"=\sqrt{13.18\times 5}\\

 =\sqrt{65.9}=8.11"